OpenStudy (amtran_bus):

How does 1/B-1/A = AB/A-B?

OpenStudy (amtran_bus):

@dan815 @DLS

OpenStudy (anonymous):

\[\frac{ 1 }{ \frac{ 1 }{ b } -\frac{ 1 }{ a }}=\frac{ ab }{ a-b }\]

OpenStudy (amtran_bus):

But why do you put them over 1?

OpenStudy (anonymous):

that one you posted is incorrect\[\frac{ 1 }{ b }-\frac{ 1 }{ a}=\frac{ a-b }{ ab } \]

OpenStudy (amtran_bus):

Thank you! Missed that!

OpenStudy (amtran_bus):

@Jonask , why did you put it all over1?

OpenStudy (anonymous):

well to make it equivalent to your result otherwise it wont work if you just leave it like that its not true \[\huge \frac{ a }{ b }+\frac{ c }{ d }=\frac{ ad+bc }{ bd }\]

OpenStudy (amtran_bus):

Will you please also show me the steps for 1/b-1/a = a-b/ab? I need to see how you got it?

OpenStudy (amtran_bus):

Thanks so much.

OpenStudy (amtran_bus):

Nvm, you just did. God bless.